A square and rectangle have the same area. The length of the rectangle is five inches more than twice the length of the side of the square. The width of the rectangle is 6 inches less than the side of the square. Find the length of the side of the square.
Answer by Boreal(15235)

square has side x
rectangle has length 2x+5 and width x-6. Its area is 2x^2-7x-30
The square's area is x^2
The difference between the two is zero, since they are equal
so x^2-7x-30=0
(x-10)(x+3)=0
x=10 inches side of square
rectangle is 25 x 4 inches